OpenStudy (anonymous):

Vivian combined different amounts of white paint, blue paint, and green paint to make 145 milliliters (ml) of paint of a new shade. The amount of white paint was x ml. The amount of blue paint was 10 ml less than twice the amount of white paint, and the amount of green paint was half the amount of blue paint. This is represented in the equation below. x + (2x − 10) + 1/2(2x − 10) = 145 What was the difference in the amounts of white paint and blue paint Vivian combined?

OpenStudy (anonymous):

If it helps this are the possible answers: 30 ml 35 ml 40 ml 45 ml

OpenStudy (anonymous):

first solve for x, that will give you amount of white paint x + 2x - 10 + x - 5 = 145 4x = 145 + 15 x = 150/4 = 37.5 ml now you can find the amounts for the rest: blue: 2x-10 = 2(37.5) - 10 = 65 ml green: 1/2(blue) = 65/2 = 32.5 ml

OpenStudy (anonymous):

So now you subtract 65-37.5?

OpenStudy (anonymous):

sorry, x = 40

OpenStudy (anonymous):

whilte paint = 40 ml blue: 2(40)-10 = 70 different of blue and white 70-40 = 30 ml

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Is that from flvs? cause im taking the course too :)